OpenStudy (anonymous):

First and second derivatives of (x+1)^(x+1) . i've tried wolfram but didn't understand the steps ... specially for the second derivative

OpenStudy (anonymous):

isn't (x+1)^(x+1)=(x+1)(x+1)^x or am i wrong ?

OpenStudy (zarkon):

you can't do that

OpenStudy (zarkon):

use \[f(x)=e^{\ln(f(x))}\]

OpenStudy (anonymous):

(x+1)^(x+1)=e^[(x+1)ln(x+1)]

OpenStudy (anonymous):

i don't know that rule .. where can i found the proof of it ?

OpenStudy (zarkon):

e and ln are inverse functions

OpenStudy (zarkon):

\[f^{-1}(f(g(x))=g(x)\]

OpenStudy (anonymous1):

To find the first derivative, use logarithmic differentiation. Take the expression: \[y = (x+1)^{x+1} \] and take the logarithm of both sides: \[\ln |y| = \ln |(x+1)^{x+1}| \] Now use implicit differentiation to find \[\frac{dy}{dx}\]which is the desired result. For that, use the following rule: \[\frac{d}{dx}(\ln|x|)=\frac{1}{x}\]

OpenStudy (anonymous1):

Notice that \[\ln|y|\] is the natural logarithm of the absolute value of y (between two vertical bars: |y|).